# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ Abstract
This paper investigates signal estimation in wireless transmission from the perspective of statistical machine learning, where the transmitted signals may be from an integrated sensing and communication system; that is, 1) signals may be not only discrete constellation points but also arbitrary complex values; 2) signals may be spatially correlated. Particular attention is paid to handling various uncertainties such as the uncertainty of the transmitting signal covariance, the uncertainty of the channel matrix, the uncertainty of the channel noise covariance, the existence of channel impulse noises (i.e., outliers), and the limited sample size of pilots. To proceed, a distributionally robust machine learning framework that is insensitive to the above uncertainties is proposed for beamforming (at the receiver) and estimation of wireless signals, which reveals that channel estimation is not a necessary operation. For optimal linear estimation, the proposed framework includes several existing beamformers as special cases such as diagonal loading and eigenvalue thresholding. For optimal nonlinear estimation, estimators are limited in reproducing kernel Hilbert spaces and neural network function spaces, and corresponding uncertainty-aware solutions (e.g., kernelized diagonal loading) are derived. In addition, we prove that the ridge and kernel ridge regression methods in machine learning are distributionally robust against diagonal perturbation in feature covariance.
@@
Wireless Transmission, Smart Antenna, Machine Learning, Robust Estimation, Robust Beamforming, Distributional Uncertainty, Channel Uncertainty, Limited Sample.
@@ I Introduction
Different from traditional paradigms, in emerging communication systems, e.g., integrated sensing and communication (ISAC) systems, transmitted signals may be arbitrary complex-valued symbols and spatially correlated [21, 22, 23]. As a result, mean-squared error is a preferred performance measure to investigate the beamforming and estimation problem of wireless signals, which is, therefore, the focus of this paper.
@@
    What is the relation between the signal-model-based approaches (e.g., Wiener and Capon receivers) and the data-driven approaches (e.g., deep-learning receivers)? In other words, how can we build a mathematically unified modeling framework to interpret all the existing digital beamformers?
@@
    In addition to the limited pilot size and the uncertainty in steering vectors, there exist other uncertainties in the signal model: the uncertainty in the transmitting signal covariance, the uncertainty of the communication channel matrix, the uncertainty of the channel noise covariance, and the presence of channel impulse noises (i.e., outliers). Therefore, how can we handle all these types of uncertainties in a unified solution framework?
@@
To this end, this paper designs a unified modeling and solution framework for beamforming and estimation of wireless signals, in consideration of the scarcity of the pilot data and the different uncertainties in the signal model.
@@ I-A Contributions
The contributions of this paper can be summarized from the aspects of machine learning theory and wireless transmission theory.
In terms of machine learning theory, we give a theoretical justification and interpretation of the popular ridge regression and kernel ridge regression (i.e., quadratic loss function plus squared-$F$-norm regularization) from the perspective of distributional robustness; see Theorems 2 and 3 (including the clarifications underneath them). To the knowledge of the authors, this is the first time to rigorously prove that the ridge and kernel ridge regression methods are distributionally robust, which enriches the theory of trustworthy machine learning.
@@
    We build a fundamentally theoretical framework for the beamforming and estimation of wireless signals from the perspective of statistical machine learning. In addition to the linear estimation method (i.e., beamforming at the receiver), nonlinear approaches (i.e., nonlinear beamforming) are also discussed in reproducing kernel Hilbert spaces and neural network function spaces. In particular, we reveal that channel estimation is not a necessary operation in beamforming and estimation of wireless signals. For details, see Subsection III-A.
@@
    The presented framework is particularly developed from the perspective of distributional robustness which can therefore combat the limited size of the training data set (i.e., limited pilot length) and several types of uncertainties in the wireless signal model such as the uncertainty in the transmitting power matrix, the uncertainty in the communication channel matrix, the existence of channel impulse noises (i.e., outliers), the uncertainty in the covariance matrix of channel noises, etc. For details, see Subsections III-B and III-C, and the technical developments in Sections IV and V.
@@
    Existing methods such as diagonal loading and eigenvalue thresholding are proven to be distributionally robust against the limited pilot size and all the aforementioned uncertainties in the wireless signal model. Extensions of diagonal loading and eigenvalue thresholding are proposed as well. Moreover, the kernelized diagonal loading and the kernelized eigenvalue thresholding methods are put forward for nonlinear estimation cases. For details, see Examples 1-7 and Subsections IV-B.
@@
    The distributionally robust beamforming and estimation problems across multiple frames, where channel conditions may change, are also investigated. For details, see Subsections IV-C and V-A2.
@@ I-B Notations
be the real-space representations of $\bm{H}$ where $\underline{\bm{H}}\in\mathbb{R}^{2N\times M}$ and $\underline{\underline{\bm{H}}}\in\mathbb{R}^{2N\times 2M}$. The running index set induced by an integer $N$ is defined as $[N]\coloneqq\{1,2,\ldots,N\}$. To concatenate matrices and vectors, MATLAB notations are used: i.e., $[\bm{A},~{}\bm{B}]$ for row stacking and $[\bm{A};~{}\bm{B}]$ for column stacking. We let $\bm{\Gamma}_{M}\coloneqq[\bm{I}_{M},~{}\bm{J}_{M}]\in\mathbb{C}^{M\times 2M}$ where $\bm{I}_{M}$ denotes the $M$-dimensional identity matrix, $\bm{J}_{M}\coloneqq j\cdot\bm{I}_{M}$, and $j$ denotes the imaginary unit. Let $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ denote a real Gaussian distribution with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$. We use $\mathcal{CN}(\bm{s},\bm{P},\bm{C})$ to denote a complex Gaussian distribution with mean $\bm{s}$, covariance $\bm{P}$, and pseudo-covariance $\bm{C}$; if $\bm{C}$ is not specified, we imply $\bm{C}=\bm{0}$. Throughout the paper, we use $\bm{R}_{xs}\coloneqq\mathbb{E}\mathbf{x}\mathbf{s}^{\mathsf{H}}$ to denote the covariance matrix of $\mathbf{x}$ and $\mathbf{s}$.
@@ III Problem Formulation
where $\mathbf{x}\in\mathbb{C}^{N}$ is the received signal, $\mathbf{s}\in\mathbb{C}^{M}$ is the transmitted signal, $\bm{H}\in\mathbb{C}^{N\times M}$ is the channel matrix, and $\mathbf{v}\in\mathbb{C}^{N}$ is the zero-mean channel noise. The precoding operation (if exists) is integrated in $\bm{H}$. We assume $\mathbf{x}\sim\mathcal{CN}(\bm{0},\bm{R}_{x},\bm{C}_{x})$, $\mathbf{s}\sim\mathcal{CN}(\bm{0},\bm{R}_{s},\bm{C}_{s})$, and $\mathbf{v}\sim\mathcal{CN}(\bm{0},\bm{R}_{v})$. The transmitted symbols $\mathbf{s}$ have zero means, which may be not only discrete values from constellations such as quadrature amplitude modulation but also arbitrary symbols such as integrated sensing and communication (ISAC) signals. We consider $L$ pilots/snapshots $\bm{S}\coloneqq(\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L})$ in each frame, and the corresponding received symbols are $\bm{X}\coloneqq(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L})$ under the noise sequence $(\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{L})$; in each frame, $\bm{H}$ is assumed to be unchanged. For generality, we suppose that the covariances of the signal $\mathbf{s}$ and the noise $\mathbf{v}$ are $\bm{R}_{s}\coloneqq\bm{P}\succeq\bm{0}$ and $\bm{R}_{v}\coloneqq\bm{R}\succeq\bm{0}$, respectively. Note that $\bm{P}$ and $\bm{R}$ may not be identity or diagonal matrices: i.e., $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ can be correlated (e.g., in ISAC), so can be $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, for every $i,j\in[L]$. Consider the real-space representation of the signal model (5) by stacking the real and imaginary components:
@@ III-A1 Optimal Nonlinear Estimation (Nonlinear Beamforming)
$\hat{\mathbf{s}}=\bm{\phi}(\mathbf{x})=\mathbb{E}({\mathbf{s}|\mathbf{x}}),$ (8)
@@
where $\mathbf{e}$ denotes the signal’s estimation error; (9) implies that the estimation problem of $\mathbf{s}$ from $\mathbf{x}$ can be seen as a nonlinear regression problem of the collected data pairs $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$. In literature, some authors refer to nonlinear estimator $\bm{\phi}(\cdot)$ as a nonlinear beamformer. Usually, it is technically complicated to find the optimal $\bm{\phi}(\cdot)$ from the whole space $\mathcal{B}$ of Borel-measurable functions. Therefore, in practice, we may find the optimal approximation of $\bm{\phi}(\cdot)$ in an RKHS $\mathcal{H}$ or a NNFS $\mathcal{K}$; note that $\mathcal{H}$ and $\mathcal{K}$ are two subspaces of $\mathcal{B}$. However, both $\mathcal{H}$ and $\mathcal{K}$ are sufficiently rich in the sense of being able to include all continuous bounded functions.
@@ III-A2 Optimal Linear Estimation (Beamforming)
where $\bm{W}\in\mathbb{C}^{M\times N}$ is called a smart antenna or a beamformer. In this linear case, (7) reduces to the usual Wiener–Hopf beamforming problem
@@
where $\bm{R}_{x}\coloneqq\mathbb{E}{\mathbf{x}\mathbf{x}^{\mathsf{H}}}\in\mathbb{C}^{N\times N}$, $\bm{R}_{s}\coloneqq\mathbb{E}{\mathbf{s}\mathbf{s}^{\mathsf{H}}}=\bm{P}\in\mathbb{C}^{M\times M}$, and $\bm{R}_{xs}\coloneqq\mathbb{E}{\mathbf{x}\mathbf{s}^{\mathsf{H}}}\in\mathbb{C}^{N\times M}$. Denote $\bm{R}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{H}}=\bm{R}\in\mathbb{C}^{N\times N}$. Since $\bm{R}_{x}=\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}=\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R}$ and $\bm{R}_{xs}=\bm{H}\bm{R}_{s}+\bm{R}_{sv}=\bm{H}\bm{P}$, the solution of (12), or (11), is
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{Wiener}}&=\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\\ &=\bm{P}\bm{H}^{\mathsf{H}}[\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R}]^{-1},\end{array}$ (13)
@@ III-A3 Role of Channel Estimation
As we can see from (7) and (11), channel estimation is not a necessary step in the estimation of wireless signals. The only necessary element, from the perspectives of statistical inference and machine learning, is the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ of the received signal $\mathbf{x}$ and the transmitted signal $\mathbf{s}$. Therefore, the following two points can be highlighted.
@@
Since $\bm{H}$ can be an integrated channel after precoding (e.g., $\bm{H}\coloneqq\bm{H}_{0}\bm{B}_{p}$ where $\bm{H}_{0}$ denotes the physical channel and $\bm{B}_{p}$ the precoder), the knowledge of the precoding operation is also unnecessary for receivers when estimating the transmitted signals. This is practically beneficial because we do not need to build an exact channel model and then accurately estimate the channel, which is challenging especially when the channel statistics are time-varying or the channel noise is non-Gaussian and non-stationary; cf. [10].
III-B Issue of Distributional Uncertainty
In practice, the true joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is unknown but we have samples $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$ from $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. Hence, the beamforming and estimation problem of wireless signals is a data-driven statistical inference (i.e., statistical machine learning) problem. We let
@@ III-B Issue of Distributional Uncertainty
denote the empirical distribution supported on the $L$ collected data $\{(\bm{x}_{i},\bm{s}_{i})\}_{i\in[L]}$, where $\delta_{(\bm{x}_{i},\bm{s}_{i})}$ denotes the Dirac distribution (i.e., point-mass distribution) centered on $(\bm{x}_{i},\bm{s}_{i})$; note that ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ is a discrete distribution. If we use the estimated joint distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ as a surrogate for the true joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$, (7) becomes the conventional empirical risk minimization (ERM) $\min_{\bm{\phi}\in\mathcal{B}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}$, i.e.,
@@
There exists the distributional difference between the sample-defined nominal distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and true data-generating distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ due to the limited size of the training data set (i.e., limited pilot length), the uncertainty in the transmitting power matrix, the uncertainty in the channel matrix (e.g., the channel may be time-selective in a frame), the existence of channel impulse noises (i.e., outliers), the uncertainty in channel noise covariance, etc. From the perspective of applied statistics and machine learning, the distributional difference between ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ (i.e., the distributional uncertainty of ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ compared to $\mathbb{P}_{\mathbf{x},\mathbf{s}}$) may cause significant performance degradation of (15) compared to (7), so is the performance deterioration of (16) compared to (12). For extensive reading on this point, see Appendix C of the online supplementary materials. Therefore, to reduce the adverse effect introduced by the distributional uncertainty in ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$, a new surrogate of (7) rather than the sample-averaged approximation in (15) is expected.
III-C Distibutionally Robust Estimation

### v2

# Distributionally Robust Receive Beamforming

@@ Abstract
This article investigates signal estimation in wireless transmission (i.e., receive beamforming) from the perspective of statistical machine learning, where the transmit signals may be from an integrated sensing and communication system; that is, 1) signals may be not only discrete constellation points but also arbitrary complex values; 2) signals may be spatially correlated. Particular attention is paid to handling various uncertainties such as the uncertainty of the transmit signal covariance, the uncertainty of the channel matrix, the uncertainty of the channel noise covariance, the existence of channel impulse noises, and the limited sample size of pilots. To proceed, a distributionally robust machine learning framework that is insensitive to the above uncertainties is proposed, which reveals that channel estimation is not a necessary operation. For optimal linear estimation, the proposed framework includes several existing beamformers as special cases such as diagonal loading and eigenvalue thresholding. For optimal nonlinear estimation, estimators are limited in reproducing kernel Hilbert spaces and neural network function spaces, and corresponding uncertainty-aware solutions (e.g., kernelized diagonal loading) are derived. In addition, we prove that the ridge and kernel ridge regression methods in machine learning are distributionally robust against diagonal perturbation in feature covariance.
@@
Wireless Transmission, Smart Antenna, Machine Learning, Robust Estimation, Robust Beamforming, Distributional Uncertainty, Channel Uncertainty, Limited Pilot.
@@ I Introduction
Different from traditional paradigms, in emerging communication systems, e.g., integrated sensing and communication (ISAC) systems, transmitted signals may be arbitrary complex-valued symbols and spatially correlated [21, 22, 23]. As a result, mean-squared error is a preferred performance measure to investigate the receive beamforming and estimation problem of wireless signals, which is, therefore, the focus of this article.
@@
    What is the relation between the signal-model-based approaches (e.g., Wiener and Capon receivers) and the data-driven approaches (e.g., deep-learning receivers)? In other words, how can we build a mathematically unified modeling framework to interpret all the existing digital receive beamformers?
@@
    In addition to the limited pilot size and the uncertainty in steering vectors, there exist other uncertainties in the signal model: the uncertainty in the transmit signal covariance, the uncertainty of the communication channel matrix, the uncertainty of the channel noise covariance, and the presence of channel impulse noises (i.e., outliers). Therefore, how can we handle all these types of uncertainties in a unified solution framework?
@@
To this end, this article designs a unified modeling and solution framework for receive beamforming of wireless signals, in consideration of the scarcity of the pilot data and the different uncertainties in the signal model.
@@ I-A Contributions
The contributions of this article can be summarized from the aspects of machine learning theory and wireless transmission theory.
In terms of machine learning theory, we give a justification of the popular ridge regression and kernel ridge regression (i.e., quadratic loss function plus squared-$F$-norm regularization) from the perspective of distributional robustness against diagonal perturbation in feature covariance, which enriches the theory of trustworthy machine learning; see Theorems 2 and 3, as well as Corollaries 3 and 5.
@@
    We build a fundamentally theoretical framework for receive beamforming from the perspective of statistical machine learning. In addition to the linear estimation methods, nonlinear approaches (i.e., nonlinear beamforming) are also discussed in reproducing kernel Hilbert spaces and neural network function spaces. In particular, we reveal that channel estimation is not a necessary operation in receive beamforming. For details, see Subsection III-A.
@@
    The presented framework is particularly developed from the perspective of distributional robustness which can therefore combat the limited size of pilot data and several types of uncertainties in the wireless signal model such as the uncertainty in the transmit power matrix, the uncertainty in the communication channel matrix, the existence of channel impulse noises (i.e., outliers), the uncertainty in the covariance matrix of channel noises, etc. For details, see Subsection III-B, and the technical developments in Sections IV and V.
@@
    Existing methods such as diagonal loading and eigenvalue thresholding are proven to be distributionally robust against the limited pilot size and all the aforementioned uncertainties in the wireless signal model. Extensions of diagonal loading and eigenvalue thresholding are proposed as well. Moreover, the kernelized diagonal loading and the kernelized eigenvalue thresholding methods are put forward for nonlinear estimation cases. For details, see Corollary 1, Examples 4 and 5, and Subsections IV-B.
@@
    The distributionally robust receive beamforming and signal estimation problems across multiple frames, where channel conditions may change, are also investigated. For details, see Subsections IV-C and V-A2.
@@ I-B Notations
be the real-space representations of $\bm{H}$ where $\underline{\bm{H}}\in\mathbb{R}^{2N\times M}$ and $\underline{\underline{\bm{H}}}\in\mathbb{R}^{2N\times 2M}$. The running index set induced by an integer $N$ is defined as $[N]\coloneqq\{1,2,\ldots,N\}$. To concatenate matrices and vectors, MATLAB notations are used: i.e., $[\bm{A},~{}\bm{B}]$ for row stacking and $[\bm{A};~{}\bm{B}]$ for column stacking. We let $\bm{\Gamma}_{M}\coloneqq[\bm{I}_{M},~{}\bm{J}_{M}]\in\mathbb{C}^{M\times 2M}$ where $\bm{I}_{M}$ denotes the $M$-dimensional identity matrix, $\bm{J}_{M}\coloneqq j\cdot\bm{I}_{M}$, and $j$ denotes the imaginary unit. Let $\mathcal{N}(\bm{\mu},\bm{\Sigma})$ denote a real Gaussian distribution with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$. We use $\mathcal{CN}(\bm{s},\bm{P},\bm{C})$ to denote a complex Gaussian distribution with mean $\bm{s}$, covariance $\bm{P}$, and pseudo-covariance $\bm{C}$; if $\bm{C}$ is not specified, we imply $\bm{C}=\bm{0}$.
@@ III Problem Formulation
where $\mathbf{x}\in\mathbb{C}^{N}$ is the received signal, $\mathbf{s}\in\mathbb{C}^{M}$ is the transmitted signal, $\bm{H}\in\mathbb{C}^{N\times M}$ is the channel matrix, and $\mathbf{v}\in\mathbb{C}^{N}$ is the zero-mean channel noise. The precoding operation (if exists) is integrated in $\bm{H}$. The transmitted symbols $\mathbf{s}$ have zero means, which may be not only discrete values from constellations such as quadrature amplitude modulation but also arbitrary symbols such as integrated sensing and communication (ISAC) signals. We consider $L$ pilots $\bm{S}\coloneqq(\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L})$ in each frame, and the corresponding received symbols are $\bm{X}\coloneqq(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L})$ under the noise $(\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{L})$. We suppose that $\bm{R}_{s}\coloneqq\mathbb{E}\mathbf{s}\mathbf{s}^{\mathsf{H}}$ and $\bm{R}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{H}}$ may not be identity or diagonal matrices: i.e., the components of $\mathbf{s}$ can be correlated (e.g., in ISAC), so can be these of $\mathbf{v}$. Consider the real-space representation of the signal model (5) by stacking the real and imaginary components:
@@ III-A1 Optimal Nonlinear Estimation (Nonlinear Beamforming)
$\hat{\mathbf{s}}=\bm{\phi}(\mathbf{x})=\mathbb{E}({\mathbf{s}|\mathbf{x}}).$ (8)
@@
where $\mathbf{e}$ denotes the signal’s estimation error; (9) implies that, in practice where $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is unknown, the data-driven estimation problem of $\mathbf{s}$ given $\mathbf{x}$ can be seen as a nonlinear regression problem of the collected data pairs $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$. Usually, it is technically complicated to find the optimal $\bm{\phi}(\cdot)$ from the whole space $\mathcal{B}$ of Borel-measurable functions. Therefore, in practice, we may find the optimal approximation of $\bm{\phi}(\cdot)$ in an RKHS $\mathcal{H}$ or a NNFS $\mathcal{K}$; note that $\mathcal{H}$ and $\mathcal{K}$ are two subspaces of $\mathcal{B}$. However, both $\mathcal{H}$ and $\mathcal{K}$ are sufficiently rich because they can be dense in the space of all continuous bounded functions.
@@ III-A2 Optimal Linear Estimation (Beamforming)
where $\bm{W}\in\mathbb{C}^{M\times N}$ is called a smart antenna or a receive beamformer. In this linear case, (7) reduces to the usual Wiener–Hopf beamforming problem
@@
where $\bm{R}_{x}\coloneqq\mathbb{E}{\mathbf{x}\mathbf{x}^{\mathsf{H}}}\in\mathbb{C}^{N\times N}$ and $\bm{R}_{xs}\coloneqq\mathbb{E}{\mathbf{x}\mathbf{s}^{\mathsf{H}}}\in\mathbb{C}^{N\times M}$. Since $\bm{R}_{x}=\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}$ and $\bm{R}_{xs}=\bm{H}\bm{R}_{s}+\mathbb{E}{\mathbf{v}\mathbf{s}^{\mathsf{H}}}=\bm{H}\bm{R}_{s}$, the solution of (12), or (11), is
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{Wiener}}&=\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\\ &=\bm{R}_{s}\bm{H}^{\mathsf{H}}[\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}]^{-1},\end{array}$ (13)
@@ III-A3 Role of Channel Estimation
Eqs. (7) and (11) imply that channel estimation is not a necessary step in receive beamforming. The only necessary element, from the perspective of statistical machine learning, is the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ of the received signal $\mathbf{x}$ and the transmitted signal $\mathbf{s}$. Therefore, the following two points can be highlighted.
@@
III-B Distributional Uncertainty and Distributional Robustness
For ease of conceptual illustration, we start with the following stationary-channel assumption in this subsection: The channel statistics remain unchanged within the communication frame so that the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is fixed over time. That is, pilot data $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$ and non-pilot communication data are drawn from the same unknown distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. For the general case where the channel is not statistically stationary within a frame, see Appendix C of the online supplementary materials; the statistical non-stationarity of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ may be due to the time-selectivity of the transmit power matrix $\bm{R}_{s}$, of the channel matrix $\bm{H}$, and/or of the channel noise covariance $\bm{R}_{v}$.
III-B1 Issue of Distributional Uncertainty
In practice, the true joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is unknown but can be estimated by the pilot data. Hence, the estimation of wireless signals is a data-driven statistical inference (i.e., statistical machine learning) problem. We let
@@ III-B1 Issue of Distributional Uncertainty
denote the empirical distribution supported on the $L$ collected data $\{(\bm{x}_{i},\bm{s}_{i})\}_{i\in[L]}$, where $\delta_{(\bm{x}_{i},\bm{s}_{i})}$ denotes the Dirac distribution (i.e., point-mass distribution) centered on $(\bm{x}_{i},\bm{s}_{i})$; note that ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ is a discrete distribution. If we use the estimated joint distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ as a surrogate of the true joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$, (7) becomes the conventional empirical risk minimization (ERM) $\min_{\bm{\phi}\in\mathcal{B}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}$, i.e.,
@@
There exists the distributional difference between the sample-defined nominal distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and true data-generating distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ due to the limited size of the training data set (i.e., limited pilot length) and the time-selectivity of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. From the perspective of applied statistics and machine learning, the distributional difference between ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ (i.e., the distributional uncertainty of ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ compared to $\mathbb{P}_{\mathbf{x},\mathbf{s}}$) may cause significant performance degradation of (15) compared to (7), so is the performance deterioration of (16) compared to (12). For extensive reading on this point, see Appendix C of the online supplementary materials. Therefore, to reduce the adverse effect introduced by the distributional uncertainty in ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$, a new surrogate of (7) rather than the sample-averaged approximation in (15) is expected.
III-B2 Distibutionally Robust Estimation